OpenStudy (anonymous):

To which graph does the point (8, −2) belong? y ≤ −x + 4 y ≤ x − 11 y ≤ −2x − 3 y ≤ 2x + 1

OpenStudy (zzr0ck3r):

let x=8 and y = -2 which one makes since when you plug those in?

OpenStudy (anonymous):

B? @zzr0ck3r

OpenStudy (anonymous):

@zzr0ck3r b?

OpenStudy (anonymous):

@wio

OpenStudy (zzr0ck3r):

\[y\le-x+4\]let x = 8 and y = -2 \[-2\le-8+4=-4\]this does not make sense for a \[y\le x-11\]let x = 8 and y = -2 \[-2\le 8-11=-3\]this does not make sense \[y\le -2x-3\]let x = 8 and y = -2\[\]\[-2\le -2(8)-3=-19\]this does not make sense \[y\le2x+1\]let x = 8 and y = -2\[-2\le 2(8)+1=17 \\\color{red}{\text{DING DING DING}!}\] this makes sense.

OpenStudy (zzr0ck3r):

@Nurialozza96 understand?

OpenStudy (anonymous):

Def. I appreciate that you wrote down step by step ! thankyou